# https://www.intelligence.tuc.gr/gogs/

### v2

@@ 3 Deep Reinforcement Learning
RL [bertsekas2019reinforcement, Kaelbling1996Reinforcement, Sutton1998Reinforcement] is a specialized area within AI. It equips an agent with the ability to learn in an environment by repeatedly interacting with it autonomously; and, importantly, it allows for the optimization of the agent’s sequential behavior under uncertainty. In more detail, the agent observes the representation of the current state of an environment, takes an action that leads to a new state, possibly obtains a reward, observes the new current state, and so on. The agent’s ultimate goal is to learn a policy, i.e. a mapping of environment states to actions, one which would maximize its cumulative reward.
RL is naturally paired with the Markov Decision Processes (MDPs) framework [howard1960dynamic, Puterman1994MarkovDP, Wal1985Markov], which provides a more detailed and rigorous portrayal of an evolving environment and its dynamics—albeit under certain assumptions, such as stationrity, that might or might not apply in the real world. MDPs are employed for modeling decision making problems, and are an extension of Markov Chains [Joseph2020Markov, Sargent1960Finite]. By definition, the MDP is a tuple $(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma)$ where:
@@
  • $\mathcal{T}(s,\mathrm{a},s^{\prime})$ = $\mathcal{P}_{r}(s^{\prime}|s,\mathrm{a})$ is the transition function - a distribution that returns the probability of landing to state $s^{\prime}$ by doing action $\mathrm{a}$ in state $s$
  • $\mathcal{R}(s,\mathrm{a})$ is the reward function
@@
Noticeably, the transition probabilities and the reward function depend only on the current state $s$ and not on all past states (Markov property). Another assumption in this formulation, is that the agent has full observability over the environmental state. However, this is not always the case since in many problems the agent can only observe certain parts of the underlying hidden state in effect. To this end, an extension to MDP has been presented, named POMDP [kaelbling1998planning, monahan1982state, Spaan2005Perseus:] that addresses partial observability. A POMDP is defined as a tuple $(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\Omega,\mathcal{O},\gamma)$ where:
@@
  • $\mathcal{O}(s^{\prime},\mathrm{a},o)=P_{r}(o|s^{\prime},a)$ is the observation function that generates the conditional probability for receiving observation $o$ after taking action $\mathrm{a}$ and landing in state $s^{\prime}$
A single interaction with the environment goes as follows: The agent lies in a state $s\in\mathcal{S}$, decides to take an action $\mathrm{a}\in\mathcal{A}$ which will result to a new state $s^{{}^{\prime}}\in\mathcal{S}$ with probability $\mathcal{T}(s,\mathrm{a},s^{\prime})$ and receives an immediate reward $\mathcal{R}(s,\mathrm{a})$ as well as an observation $o\in\Omega$ with probability $\mathrm{O}(s^{\prime},\mathrm{a},o)$, if partial information is considered. The discount factor $\gamma$ also plays an important role in the learning competence of the algorithm, as it determines how much are immediate rewards favored over future rewards. If $\gamma=0$ the agent becomes "myopic", taking only the actions that will maximize its immediate reward. As $\gamma$ approaches 1, the agent becomes more far-sighted, taking actions that maximize its long term reward. Typically, $\gamma$ is kept constant throughout the whole training phase and its exact value is selected based on the characteristics of the task at hand.
@@
$G_{t}=E(\sum_{t=0}^{\infty}\gamma^{t}\mathcal{R}(s_{t},\mathrm{a}_{t}))$
@@
As the agent keeps interacting with the environment, it collects data in the form of trajectories of actions, observations and rewards. These experiences are utilized by the RL algorithm in order to update its current policy $\pi$ and make better decisions in the future. Under these circumstances, there are two main approaches to addressing decision-making challenges in the context of RL; model-based methods and model-free methods [Nagabandi2017Neural, Xu2018Algorithmic].
@@
Model-free methods, on the other hand, do not require a model of the environment to operate and rely solely on the experience gathered from the agent interacting with the environment. In turn, this family of algorithms can be classified into two frameworks; value-based algorithms and policy-based algorithms [Arulkumaran2017Deep, Liu2020Overview]. In value-based algorithms, the agent maintains a value function (usually in the form of a table) that maps state-action pairs ($s$, $\mathrm{a}$) (or just states) to a single, real, value. One of the most well-known algorithms in this category - and generally in RL - is $Q$-learning [watkins1992q]. $Q$-learning, maintains a $Q$-table with values $Q(s,\mathrm{a})$. These $Q$-values serve as an estimate of the expected utility of taking action $\mathrm{a}$ in state $s$. This method dictates the agent to iteratively update its $Q$-values following the Bellman equation:
@@
On the other hand, policy-based RL algorithms focus on maximizing directly the parameterized policy’s rewards, through adjusting its parameters, instead of evaluating states. These methods often utilize an optimization algorithm to update the parameters efficiently, such as gradient ascent [Brereton2009Steepest], and therefore known more specifically as policy-gradient algorithms. One classic algorithm of this family is REINFORCE [Williams1992Simple]. REINFORCE updates the policy parameters based on the objective function $J(\theta)$:
@@
Actor-critic RL algorithms combine the strengths of value-based and policy-based methods, offering a balanced approach for solving RL problems. In these algorithms, the “actor” updates the policy based on feedback from the “critic”, which offers value estimation of the current policy. This synergy of different paradigms allows for more robust and efficient learning, and mitigates unilateral problems such as high variance and slow convergence. In Figure 4 we can see a visualized taxonomy of the different families of RL algorithms, as presented in [canese2021multi].
Figure 4: A taxonomy of RL algorithms [canese2021multi]. Note that there exist many algorithms that are either policy-based or value-based and are not actor-critic.
@@
Typically, we aim for our agent to explore more in the early stages of training and exploit more as training progresses. So it seems only natural to anneal the $\epsilon$ probability from a high initial value (usually 1) to a low value (usually $\leq 0.1$) as training progresses. More sophisticated exploration strategies exist with potentially better performance than $\epsilon$-greedy, such as Boltzmann (also referred to as Softmax) Exploration [cesa2017boltzmann], Upper Confidence Bound [auer2002finite] and Thompson Sampling [Russo2017A].
@@
This serious challenge can be partially overcome with the use of non-linear function approximation [Goddard1965Approximation] NN-based techniques, such as DL methods [Deng2014Deep, LeCun2015Deep]. DL involves deep NNs with many hidden layers, typically ranging from at least one to often two to ten hidden layers. This structure makes the model able to adapt to more complex functions and capture high-level features in input data. As expected, in recent years, methods using DL have excelled in intricate problems with high dimensionality such as image recognition and natural language processing, including advanced applications like large language models [Zhao2023A] and generative adversarial networks [Goodfellow2014Generative] that produce human-like text and highly realistic images.
To this end, DL and NNs have given rise to DRL enabling RL to tackle problems that were previously deemed infeasible due to the curse of dimensionality. Of course, most of the aforementioned core frameworks of classic RL, such as MDPs, POMDPs, algorithm categories, solution concepts and exploration strategies, transfer over to the domain of DRL.
In recent years, research in DRL has sparked off a wide array of methodologies that push the boundaries of the field. By combining DL with already-existing RL literature several advancements have been made. These advancements include policy-based algorithms like Proximal Policy Optimization (PPO) [schulman2017proximal] and value-based algorithms like Deep $Q$-Network (DQN) [Mnih2013Playing, Mnih2015Human-level]. Also, actor-critic models that intertwine value and policy based methods to benefit from both, have been formulated, such as Soft Actor-Critic (SAC) [haarnoja2018soft] and Deep Deterministic Policy Gradient (DDPG) [lillicrap2019continuous]. These new approaches have also led to more advanced techniques such as imitation learning, where agents can observe expert demonstrations to learn faster. Algorithms from the field of DRL have since been applied to a diverse plethora of problems, both discrete and continuous, such as robotics [Gu2016Deep, kober2013reinforcement], Atari games [Mnih2013Playing, mnih2015human], the board game of Go [Silver2016Mastering], and autonomous driving [karalakou2023deep, Sallab2017Deep, Wang2018Deep], to name a few.
@@
A lot of algorithms have stemmed from DQN in recent years with remarkable results in discrete domains. As an example, [orfanoudakis2023novel] puts forward a novel approach to managing distributed energy resources in the context of the Smart Grid by employing DQN to predict the most beneficial combinations of local flexibility estimators. Another notable work is that of [chrysomallis2023deep] which enhances DQN with implicit imitation learning in order to tackle the lane-free autonomous driving problem. The proposed algorithm makes use of state transitions, obtained from a mentor, to accelerate learning. An important spin-off of DQN is Bootstrapped DQN [osband2018randomized, osband2016deep] that employs several network heads to benefit from the diverse outputs they generate. Essentially, instead of learning a single policy, Bootstrapped DQN learns multiple policies by utilizing the inherent uncertainty of the distinct heads. At the start of each episode, a head is sampled uniformly at random, to act on the environment and collect data. The data collected can be stored in the same replay buffer, but a random distribution is used to dictate which heads are to use each experience sample. This method has been shown to outperform $\epsilon$-greedy, especially in problems that require deep exploration.
@@
Despite the achievements in the field of DRL, there are still several problems that limit its use in real-world problems. Some of the most important issues are sample inefficiency, inability to generalize to new states/problems, the demand for extensive computational resources and the obligation for a simulator to exist. Sample inefficiency refers to the requirement of vast amounts of data to even learn a substantially good policy. Even algorithms that are deemed sample-efficient are not efficient enough to be applied to many real-world tasks that cannot offer numerous samples or that even acquiring new samples is expensive. Additionally, most algorithms cannot generalize their knowledge to other domains or even to underexplored parts of the same environment. This lack of generalization means that DRL methods need to be trained again even if the environment is slightly different - a burden that can be resource exhaustive. As it happens, training a DRL algorithm, often with more than one neural networks, demands a large amount of computational resources for extended periods of time, often several days. Consequently, this issue only, can make DRL impractical for individuals that do not have access or cannot afford such computational power. Finally, many physical problems can be proven to be expensive to train an agent in, especially tasks in which the agent has a physical substance. For example, if the task is to train a real-world physical robot, the decisions made by the algorithm can damage the agent. For these tasks there need to be a simulator that can simulate both the agent and the environment in a digital form to allow for a faster and safer training procedure. Although, as is evident, such simulators are not always available and their implementation introduces an additional hurdle.
3.1 Multiagent learning
@@
An RL direction that has gained momentum over the last years is MARL [marl-book, gronauer2022multi, oroojlooy2023review]. MARL is an extension of single-agent RL, and it involves multiple agents that usually interact with each other and affect each other’s decisions, as well as each other’s perception of the environment. The relationships between the agents may vary as they may need to cooperate, compete or a mix of both behaviors, depending on the task at hand. As expected, these new dynamics that do not exist in single-agent settings, mandate the incorporation of game theory notions and techniques in the RL setting to allow for a deeper understanding of MARL and enable the analysis of strategies and outcomes. Naturally, just as DL extended the capabilities of RL, it has similarly enhanced MARL to design MADRL algorithms that are more capable in tackling complex problems. Typically, many real-world scenarios can normally be modeled with the use of multiple agents such as autonomous driving [shalev2016safe], settings with many economic entities [shavandi2022multi] and smart-grid control [Lu2020Multi-agent], to name a few. All these scenarios benefit from the ability to model multiple interconnected agents and their strategic interactions.
@@
Building on the framework of MARL, MGs [littman1994markov, littman2001value, shapley1953stochastic] (also known as stochastic games), can be considered as an extension of MDPs to contain multiple decision-making entities in a common environment. MGs formally allow for multiple agents to coexist, all of them taking actions simultaneously and encapsulate various interactions between them, be it cooperative, competitive or even a mix of both. Formally, an MG is defined as a tuple ($\mathcal{S}$, $\mathcal{A}^{i}$, $\mathcal{T}$, $\mathcal{R}^{i}$, $\gamma$) where:
@@
  • $\mathcal{A}^{i}$ is the action space for agent $i$ and $\mathcal{\mathbf{A}}=\mathcal{A}^{1}\times\mathcal{A}^{2}\times\ldots\times\mathcal{A}^{n}$ represents the joint action space
@@
  • $\mathrm{R}^{i}(s,\mathrm{a}^{i})$ is the reward function for agent $i$
@@
Despite the importance of MGs in the field, this framework does not account for partial observability in the environment from the perspective of each agent. This gap is bridged by an extension of MGs, called Partially Observable Markov Games but more commonly known as Dec-POMDPs. In the context of Dec-POMDPs, each agent operates in a subjective way, based on its individual view of the world. Considering that its point of view is likely to be incomplete, due to other agents being part of the environment, the importance of frameworks that account for partial observation, such as Dec-POMDPs, is apparent.
Formally, a Dec-POMDP can be represented as a tuple ($\mathcal{S}$, $\mathcal{A}^{i}$, $\mathcal{T}$, $\mathcal{R}^{i}$, $\Omega^{i}$, $\mathcal{O}$ $\gamma$) where:
  • $\mathcal{S}$, $\mathcal{A}^{i}$, $\mathcal{T}$, $\mathcal{R}^{i}$, $\gamma$ are the same as in MGs
  • $\Omega^{i}$ is the observation space for agent $i$ and $\mathbf{\Omega}=\Omega^{1}\times\Omega^{2}\times\ldots\times\Omega^{n}$ is the joint observation space
  • $\mathcal{O}(s^{\prime},\mathrm{a},o)=Pr(o|s^{\prime},\mathrm{a})$ is the observation function - each agent receives an observation according to $\mathcal{O}$ after making a transition in the environment
Noticeably, the differences between Dec-POMDPs and MGs are analogous to those between MDPs and POMDPs. Essentially, Dec-POMDPs can be seen as a generalization of MGs that consider partial observability of the environment. At each timestep $t$, all $n$ agents choose an action $\mathrm{a}^{i}_{t}$ and perform all the actions simultaneously. After taking that joint action $\mathrm{\mathbf{a}}_{t}$ each one receives an observation $o^{i}_{t}$ with probability $\mathcal{O}(s_{t},a^{i}_{t},o_{t})$ and an individual reward $\mathrm{r}^{i}_{t}$ from $\mathcal{R}^{i}(s_{t},\mathrm{a}^{i}_{t},s^{\prime}_{t+1})$ while the environment transits from current state $s_{t}$ to the next state $s^{\prime}_{t+1}$ with a probability corresponding to $\mathcal{T}(s_{t},\mathrm{\mathbf{a}_{t},s^{\prime}_{t+1}})$.
The solution to a Dec-POMDP is considered as a joint policy $\mathbf{\pi}^{*}$ that maximizes the expected collective sum of discounted rewards $\max_{\mathbf{\pi}}E(\sum_{t=0}^{\infty}\gamma^{t}R_{t}(s,\mathrm{\mathbf{a}}))$. Note that $R_{t}$ here is the collective reward received by all agents after the transition from timestep $t$ to timestep $t+1$. A schematic of a MARL setting is depicted in Figure 5 [marl-book].
Figure 5: A schematic of a MARL setting [marl-book]. Depicts the interactions of multiple agents within an environment
@@
Since we now maintain numerous decision-makers, it is crucial to decide how the decisions are being made from a technical perspective. There are two notable approaches, the centralized decision and the distributed decision making. In the context of the centralized decision-making, a central entity is tasked with producing joint actions $\bf{\mathrm{a}}$ based on the state of the environment. On the other hand, each agent can compute its own action based on its local observation of the state. The latter, normally enhances scalability, but often requires the use of sophisticated coordination mechanisms. As a consequence, the design of such mechanisms seems necessary, especially for cooperative scenarios where the agents are expected to reach a common goal. Another critical notion in MARL is that of agent modeling. Agent modeling focuses on designing, maintaining and updating models that are used to predict the future actions of other agents’ in the environment. This can lead to a more successful learning process since agents can anticipate the actions of other agents and adjust their strategies preemptively. The importance of agent modeling is essential in competitive scenarios (in that case it is also known as opponent modeling) where being able to forecast opponents’ moves can provide a strategic advantage.
@@
In a comprehensive analysis, [jiang2018learning] notes that communication in cooperative tasks is essential, but constant communication may impose problems in performance and energy efficiency. Consequently, they propose a model named ATOC, that aims to learn when communication is indeed necessary and how to pass information effectively. ATOC, in essence, dictates the agents to form dynamic communication groups to enhance their coordination. In more detail, this algorithm trains an attention unit with local observations and action intentions of agents to learn when communication is imperative. When the attention unit decides that communication is due, a bi-directional Long Short-Term Memory (LSTM) unit serves as a communication channel. This attention-based communication mechanism is integrated in an actor-critic model that is trained in a centralized manner, although each agent makes individual observations. In addition, this algorithm does not assume the SIH, since agents may be required to maximize the collective utility, depending on the problem at hand. The evaluation of ATOC across various cooperative problems, shows that it outperforms existing methods by enabling agents to coordinate in a more sophisticated manner.

### v3

@@ 3 Deep Reinforcement Learning
Reinforcement learning (RL) [13, 66, 133] is a specialized area within AI. It equips an agent with the ability to learn via trial and error while repeatedly interacting with an environment; and, importantly, it allows for the optimization of the agent’s sequential behavior under uncertainty. In more detail, the agent observes the representation of the current state of an environment, takes an action that leads to a new state, obtains a reward, observes the new current state, and so on. The agent’s ultimate goal is to learn a policy—i.e. a mapping of environment states to actions—that maximizes its long-term accumulated reward.
RL is naturally paired with the Markov Decision Processes (MDPs) framework [57, 107, 143], which provides a more detailed and rigorous portrayal of an evolving environment and its dynamics—albeit under certain assumptions, such as stationrity, that might or might not apply in the real world. MDPs are employed for modeling decision making problems, and are an extension of Markov Chains [65, 114]. By definition, the MDP is a tuple $(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma)$ where:
@@
  • $\mathcal{T}(s,\mathrm{a},s^{\prime})$ = $\Pr(s^{\prime}|s,\mathrm{a})$ is the transition function - a distribution that returns the probability of landing to state $s^{\prime}$ by doing action $\mathrm{a}$ in state $s$
  • $\mathcal{R}(s,\mathrm{a},s^{\prime})$ is the reward function
@@
Noticeably, the transition probabilities and the reward function depend only on the current state $s$ and not on all past states (Markov property). Another assumption in this formulation, is that the agent has full observability over the environmental state. However, this is not always the case since in many problems the agent can only observe certain parts of the underlying hidden state in effect. To this end, an extension to MDP has been presented, named POMDP [67, 91, 130] that addresses partial observability. A POMDP is defined as a tuple $(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\Omega,\mathcal{O},\gamma)$ where:
@@
  • $\mathcal{O}(s^{\prime},\mathrm{a},\omega)=\Pr(\omega|s^{\prime},a)$ is the observation function that generates the conditional probability for receiving observation $\omega\in\Omega$ after taking action $\mathrm{a}$ and landing in state $s^{\prime}$
A single interaction with the environment goes as follows: The agent lies in a state $s\in\mathcal{S}$, decides to take an action $\mathrm{a}\in\mathcal{A}$ which will result to a new state $s^{\prime}\in\mathcal{S}$ with probability $\mathcal{T}(s,\mathrm{a},s^{\prime})$ and receives an immediate reward $\mathcal{R}(s,\mathrm{a},s^{\prime})$ as well as an observation $\omega\in\Omega$ with probability $\mathrm{O}(s^{\prime},\mathrm{a},\omega)$, if partial information is considered. The discount factor $\gamma$ also plays an important role in the learning competence of the algorithm, as it determines how much are immediate rewards favored over future rewards. If $\gamma=0$ the agent becomes "myopic", taking only the actions that will maximize its immediate reward. As $\gamma$ approaches 1, the agent becomes more far-sighted, taking actions that maximize its long term reward. Typically, $\gamma$ is kept constant throughout the whole training phase and its exact value is selected based on the characteristics of the task at hand.
@@
$G_{t}=E(\sum_{t=0}^{\infty}\gamma^{t}\mathcal{R}(s_{t},\mathrm{a}_{t},s_{t+1}))$
@@
As the agent keeps interacting with the environment, it collects data in the form of trajectories of actions, observations and rewards. These experiences are utilized by the RL algorithm in order to update its current policy $\pi$ and make better decisions in the future. Under these circumstances, there are two main approaches to addressing decision-making challenges in the context of RL; model-based methods and model-free methods [94, 152].
@@
Model-free methods, on the other hand, do not require a model of the environment to operate and rely solely on the experience gathered from the agent interacting with the environment. In turn, this family of algorithms can be classified into two frameworks; value-based algorithms and policy-based algorithms [3, 81]. In value-based algorithms, the agent maintains a value function (usually in the form of a table) that maps state-action pairs ($s$, $\mathrm{a}$) (or just states) to a single, real, value. One of the most well-known algorithms in this category - and generally in RL - is $Q$-learning [146]. $Q$-learning, maintains a $Q$-table with values $Q(s,\mathrm{a})$. These $Q$-values serve as an estimate of the expected utility of taking action $\mathrm{a}$ in state $s$. This method dictates the agent to iteratively update its $Q$-values following the Bellman equation:
@@
On the other hand, policy-based RL algorithms focus on maximizing directly the parameterized policy’s rewards, through adjusting its parameters, instead of evaluating states. These methods often utilize an optimization algorithm to update the parameters efficiently, such as gradient ascent [17], and therefore known more specifically as policy-gradient algorithms. One classic algorithm of this family is REINFORCE [148]. REINFORCE updates the policy parameters based on the objective function $J(\theta)$:
@@
Actor-critic RL algorithms combine the strengths of value-based and policy-based methods, offering a balanced approach for solving RL problems. In these algorithms, the “actor” updates the policy based on feedback from the “critic”, which offers value estimation of the current policy. This synergy of different paradigms allows for more robust and efficient learning, and mitigates unilateral problems such as high variance and slow convergence. In Figure 4 we can see a visualized taxonomy of the different families of RL algorithms, as presented in [19].
Figure 4: A taxonomy of RL algorithms [19]. Note that there exist many algorithms that are either policy-based or value-based and are not actor-critic.
@@
Typically, we aim for our agent to explore more in the early stages of training and exploit more as training progresses. So it seems only natural to anneal the $\epsilon$ probability from a high initial value (usually 1) to a low value (usually $\leq 0.1$) as training progresses. More sophisticated exploration strategies exist with potentially better performance than $\epsilon$-greedy, such as Boltzmann (also referred to as Softmax) Exploration [21], Upper Confidence Bound [5] and Thompson Sampling [112].
@@
This serious challenge can be partially overcome with the use of non-linear function approximation [45] NN-based techniques, such as deep learnign (DL) methods [36, 74]. DL involves deep NNs with many hidden layers, typically ranging from at least one to often two to ten hidden layers. This structure makes the model able to adapt to more complex functions and capture high-level features in input data. As expected, in recent years, methods using DL have excelled in intricate problems with high dimensionality such as image recognition and natural language processing, including advanced applications like large language models [162] and generative adversarial networks [46] that produce human-like text and highly realistic images.
To this end, DL and NNs have given rise to deep reinforcement learning (DRL), enabling RL to tackle problems that were previously deemed infeasible due to the curse of dimensionality. Of course, most of the aforementioned core frameworks of classic RL, such as MDPs, POMDPs, algorithm categories, solution concepts and exploration strategies, transfer over to the domain of DRL.
In recent years, research in DRL has sparked off a wide array of methodologies that push the boundaries of the field. By combining DL with already-existing RL literature several advancements have been made. These advancements include policy-based algorithms like Proximal Policy Optimization (PPO) [117] and value-based algorithms like Deep $Q$-Network (DQN) [88, 89]. Also, actor-critic models that intertwine value and policy based methods to benefit from both, have been formulated, such as Soft Actor-Critic (SAC) [52] and Deep Deterministic Policy Gradient (DDPG) [76]. These new approaches have also led to more advanced techniques such as imitation learning, where agents can observe expert demonstrations to learn faster. Algorithms from the field of DRL have since been applied to a diverse plethora of problems, both discrete and continuous, such as robotics [50, 70], Atari games [88, 90], the board game of Go [126], and autonomous driving [68, 113, 145], to name a few.
@@
A lot of algorithms have stemmed from DQN in recent years with remarkable results in discrete domains. As an example, [99] puts forward a novel approach to managing distributed energy resources in the context of the Smart Grid by employing DQN to predict the most beneficial combinations of local flexibility estimators. Another notable work is that of [29], which enhances DQN with implicit imitation learning to give rise to the so-called Deep Implicit Imitation Q-Networks (DIIQN) algorithm. The approach makes use of only state transitions, obtained from a mentor, to accelerate learning, without requiring access to observations of the mentor actions; and can be applied in a host of domains, including autonomous driving. That work was later extended to allow for imitation learning when the mentor and observer agent action sets are heterogeneous, for the first time in the literature [28].
An important spin-off of DQN is Bootstrapped DQN [101, 102] that employs several network heads to benefit from the diverse outputs they generate. Essentially, instead of learning a single policy, Bootstrapped DQN learns multiple policies by utilizing the inherent uncertainty of the distinct heads. At the start of each episode, a head is sampled uniformly at random, to act on the environment and collect data. The data collected can be stored in the same replay buffer, but a random distribution is used to dictate which heads are to use each experience sample. This method has been shown to outperform $\epsilon$-greedy, especially in problems that require deep exploration.
@@
Despite the achievements in the field of DRL, there are still several problems that limit its use in real-world problems. Some of the most important issues are sample inefficiency, inability to generalize to new states/problems, the demand for extensive computational resources and the obligation for a simulator to exist. Sample inefficiency refers to the requirement of vast amounts of data to even learn a substantially good policy. Even algorithms that are deemed sample-efficient are not efficient enough to be applied to many real-world tasks that cannot offer numerous samples or that even acquiring new samples is expensive. Additionally, most algorithms cannot generalize their knowledge to other domains or even to underexplored parts of the same environment. This lack of generalization means that DRL methods need to be trained again even if the environment is slightly different - a burden that can be resource exhaustive. As it happens, training a DRL algorithm, often with more than one neural networks, demands a large amount of computational resources for extended periods of time, often several days. Consequently, this issue only, can make DRL impractical for individuals that do not have access or cannot afford such computational power. Finally, many physical problems can be proven to be expensive to train an agent in, especially tasks in which the agent has a physical substance. For example, if the task is to train a real-world physical robot, the decisions made by the algorithm can damage the agent. For these tasks there need to be a simulator that can simulate both the agent and the environment in a digital form to allow for a faster and safer training procedure. However, as is evident, such simulators are not always available and their implementation introduces an additional hurdle.
3.1 Multiagent Reinforcement learning
@@
A multiagent learning direction that has gained momentum over the last years is multiagent reinforcement learning (MARL) [1, 49, 100]. MARL is an extension of single-agent RL and it involves multiple agents that usually interact with each other and affect each other’s decisions, as well as each other’s perception of the environment. The relationships between the agents may vary as they may need to cooperate, compete or adopt a mix of both behaviors, depending on the task at hand. As expected, these new dynamics that do not exist in single-agent settings, mandate the incorporation of game theory notions and techniques in the RL setting to model these more complex MARL notions and enable the analysis of strategies and outcomes. Naturally, just as DL extends the capabilities of RL, it has similarly enhanced MARL to design multiagent DRL (MADRL) algorithms that are more effective in solving real-world problems. Typically, many real-world scenarios can normally be modeled with the use of multiple agents such as in autonomous driving [118], other settings with a multitude of economic entities [122], or smart-grid control [83], to name a few. All these scenarios benefit from the ability to model multiple interconnected agents and their strategic interactions. A schematic overview of the MARL setting is depicted in Figure 5 [1].
Figure 5: The MARL setting [1].
@@
Markov games (MGs) [1, 77, 78, 121] (also known as stochastic games), have been key for the formal definition of the MARL framework. They can be considered as an extension of MDPs to contain multiple decision-making entities in a common environment. Specifically, MGs allow for multiple co-existing agents that take actions simultaneously, and encapsulate various interactions among them—cooperative, competitive or even a mix of both. Formally, an MG is a ($\mathcal{I}$, $\mathcal{S}$, $\{\mathcal{A}^{i}\}$, $\mathcal{T}$, $\{\mathcal{R}^{i}\}$, $\gamma$, $\mu$) tuple [1], where:
  • $\mathcal{I}$ denotes the set of agents that coexist in the environment
@@
  • $\mathcal{A}^{i}$ is the action space for agent $i$ and $\mathcal{A}=\mathcal{A}^{1}\times\mathcal{A}^{2}\times\ldots\times\mathcal{A}^{n}$ represents the joint action space. A set of selected actions can then be denoted as $\mathrm{a}$
@@
  • $\mathrm{R}^{i}(s,\mathrm{a},s^{\prime})$ is the reward function for agent $i$, based on the current state $s$, the action $\mathrm{a}$, and the next state $s^{\prime}$
@@
  • $\mu:S\xrightarrow{}[0,1]$ is the initial state distribution
Despite the importance of MGs in the field, this framework does not account for partial observability in the environment from the perspective of each agent. This gap is bridged by an extension of MGs, the Partially Observable Markov Games (POMGs) [1]. In the context of POMGs, each agent operates in a subjective way, based on its individual view of the world. Considering that its point of view is likely to be incomplete due to other agents being part of the environment, the importance of frameworks that account for partial observation is apparent.
Formally, a POMG can be represented as a tuple ($\mathcal{I}$, $\mathcal{S}$, $\{\mathcal{A}^{i}\}$, $\mathcal{T}$, $\{\mathcal{R}^{i}\}$, $\{\Omega^{i}\}$, $\{\mathcal{O}^{i}\}$, $\gamma$, $\mu$) [1] where:
  • $\mathcal{I}$, $\mathcal{S}$, $\mathcal{A}^{i}$, $\mathcal{T}$, $\mathcal{R}^{i}$, $\gamma$, $\mu$ are the same as in MGs
  • $\Omega^{i}$ denotes the observation space for agent $i$
  • $\mathcal{O}^{i}(s^{\prime},\mathrm{a},\omega^{i})=\Pr(\omega^{i}|s^{\prime},\mathrm{a})$ where $\omega^{i}\in\Omega^{i}$, is the observation function; an agent $i$ receives an observation according to $\mathcal{O}^{i}$ after making a transition in the environment
In multiagent settings, at each timestep $t$, all $n$ agents choose an action $\mathrm{a}^{i}_{t}$ and perform all the actions simultaneously. After taking that joint action $\mathrm{a}_{t}$ each one receives an observation $\omega^{i}_{t+1}$ with probability $\mathcal{O}^{i}(s_{t+1},a_{t},\omega^{i}_{t+1})$ and an individual reward $\mathrm{r}^{i}_{t}$ from $\mathcal{R}^{i}(s_{t},\mathrm{a}_{t},s_{t+1})$ while the environment transitions from current state $s_{t}$ to the next state $s_{t+1}$ with a probability corresponding to $\mathcal{T}(s_{t},\mathrm{a}_{t},s_{t+1})$. At $t=0$ specifically, an initial state $s_{0}$ is sampled from $\mu$.
Noticeably, the differences between POMGs and MGs are analogous to those between MDPs and POMDPs, albeit the latter do not account for multiple agents in the environment. To this end, Decentralized POMDPs (Dec-POMDPs) ($\mathcal{I}$, $\mathcal{S}$, $\{\mathcal{A}^{i}\}$, $\mathcal{T}$, $\mathcal{R}$, $\{\Omega^{i}\}$, $\{\mathcal{O}^{i}\}$, $\gamma$, $\mu$) constitute a special case of POMGs, with the sole difference that agents share a common reward function $\mathcal{R}$ [1]:
  • $\mathcal{I}$, $\mathcal{S}$, $\mathcal{A}^{i}$, $\mathcal{T}$, $\Omega^{i}$, $\mathcal{O}^{i}$, $\gamma$, $\mu$ are the same as in POMGs
  • $\mathcal{R}$ is the global reward function, common to all agents
The absence of distinct individual reward functions dictates a cooperative aspect, as all agents seek to maximize the global rewards. The solution to a Dec-POMDP is considered as a joint policy $\mathbf{\pi}^{*}$ that maximizes the expected collective sum of discounted rewards $\max_{\mathbf{\pi}}E(\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},\mathrm{a}_{t},s_{t+1}))$. Note that $R(s_{t},\mathrm{a}_{t},s_{t+1})$ here is the global reward received by all agents after the transition from timestep $t$ to timestep $t+1$.
@@
Since we now maintain numerous decision-makers, it is crucial to decide how the decisions are being made from a technical perspective. There are two notable approaches, the centralized decision and the distributed decision making. In the context of the centralized decision-making, a central entity is tasked with producing joint actions $\bm{\mathrm{a}}$ based on the state of the environment. On the other hand, each agent can compute its own action based on its local observation of the state. The latter, normally enhances scalability, but often requires the use of sophisticated coordination mechanisms. As a consequence, the design of such mechanisms seems necessary, especially for cooperative scenarios where the agents are expected to reach a common goal. Another critical notion in MARL is that of agent modeling. Agent modeling focuses on designing, maintaining and updating models that are used to predict the future actions of other agents’ in the environment. This can lead to a more successful learning process since agents can anticipate the actions of other agents and adjust their strategies preemptively. The importance of agent modeling is essential in competitive scenarios (in that case it is also known as opponent modeling) where being able to forecast opponents’ moves can provide a strategic advantage.
@@
In a comprehensive analysis, [63] notes that communication in cooperative tasks is essential, but constant communication may impose problems in performance and energy efficiency. Consequently, they propose a model named ATOC, that aims to learn when communication is indeed necessary and how to pass information effectively. ATOC, in essence, dictates the agents to form dynamic communication groups to enhance their coordination. In more detail, this algorithm trains an attention unit with local observations and action intentions of agents to learn when communication is imperative. When the attention unit decides that communication is due, a bi-directional Long Short-Term Memory (LSTM) unit serves as a communication channel. This attention-based communication mechanism is integrated in an actor-critic model that is trained in a centralized manner, although each agent makes individual observations. In addition, this algorithm does not assume the SIH, since agents may be required to maximize the collective utility, depending on the problem at hand. The evaluation of ATOC across various cooperative problems, shows that it outperforms existing methods by enabling agents to coordinate in a more sophisticated manner.